OpenStudy (ibbutibbu.):

WILL REWARD MEDALS!!! The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank. 1/3π (x - 7)2 (x - 3) What does the factor π(x - 7)2 (x - 3) represent?

OpenStudy (imstuck):

The (x-7)^2 is the radius measure and the (x-3) is the height.

OpenStudy (ibbutibbu.):

Thank you!

OpenStudy (ibbutibbu.):

Wait...

OpenStudy (imstuck):

Volume of a cone formula is this:\[V=\frac{ 1 }{ 3 }\pi (r)^{2}h\]

OpenStudy (ibbutibbu.):

But the choices are -The volume of 3 of the same cone-shaped tanks -The area of the curved sides of the tank -The volume of 6 of the same cone-shaped tanks -The area of the base of the tank

OpenStudy (imstuck):

That would be the first choice you have there.

OpenStudy (imstuck):

The 3 from the solution gets multiplied by the 1/3 to cancel out, leaving you with only pi. Like this:

OpenStudy (imstuck):

\[3(\frac{ 1 }{ 3 })\pi r ^{2} h\]The 3's cancel out, leaving you with just\[\pi r ^{2} h\]

OpenStudy (imstuck):

get it?

OpenStudy (ibbutibbu.):

Yes, i think so. Than you!

OpenStudy (imstuck):

They tell you what the height and the radius are in the problem: height is (x-3) radius is 4 less than that: (x-3)-4=x-7 In the formula, the radius is squared and multiplied by pi and the height. See?

OpenStudy (imstuck):

Take 3 of those volume measurements and the 1/3 is gone.